# Self-assembling clusters of particles on a shrinking liquid surface

### v2

@@ IV Conclusions
In conclusion, inspired by the observation of fallen pine needles self-assembling into clusters in small puddles after rainfall, we conduct experimental investigations into the dynamic evolution of particles as the liquid surface undergoes shrinkage. We observe that particles tend to adhere to one another, forming distinct clusters as the surface area reduces. We identify the emergence of clusters and systematically explore how cluster sizes and the number of clusters evolve as a function of packing fraction. Furthermore, we analyze particle rotation during the process and discuss the formation of the final configuration comprising clusters of various orientations. Complementary numerical simulations that capture the two main factors, the adhesion between particles and the reduction of surface area, demonstrate qualitative agreement with our experimental findings. Our study contributes to a better understanding of the self-assembly phenomena occurring when macroscopic particles interact on a liquid surface whose area is shrinking.
V Materials and Methods
V.1 Preparation of particles
White PLA particles were produced via 3D printing using a Flashforge Guider II S model. Each particle measures 35 mm in length, 7 mm in width, and 7 mm in height, featuring a centrally located groove measuring 30 mm in length, 3 mm in width, and 1 mm in depth along each of its four major sides. The particles were deliberately designed to be hollow, resulting in a filling ratio of 15%, thereby improving their buoyancy on liquids. The detailed structure of the particle is illustrated in Fig. 5. Please note that the colors depicted in Fig. 5 do not represent the actual colors of the particles; the real particles are white, with grooves distinctly marked in red using a red marker to enhance visibility for the following image analysis.
@@ V.1 Preparation of particles
V.2 Image analysis
V.3 Two-particle attraction
@@
To investigate the attraction between two particles, we align them parallel to each other with their large faces facing each other. Maintaining a horizontal separation of $d=1.2$ cm, we systematically vary the initial vertical overlap length $x$ from 1/7 to 6/7 of the particle length, in increments of 1/7 as equal divisions. Upon releasing the particles, they adhere to each other, and we measure the final overlap length $y$ for each initial overlap length $x$, as depicted in Fig. 6(a). The black line represents the linear fit of the first five experimental data points, where the final overlap $y$ values are less than the particle length. The blue line denotes the particle length.
@@
Figure 6: (a) Final overlap length $y$ as a function of initial overlap length $x$ when two particles are released at a fixed horizontal distance $d=1.2$ cm. The inset illustrates the initial and final states. The black line represents the linear fit of the experimental data based on the first five data points, while the blue line denotes the particle length. (b) Determination of the adhesion of two clusters in simulations. When the outermost particles (marked in blue) of two adjacent clusters form an angle $\alpha\leq 5^{\circ}$, we rotate each cluster by $\alpha/2$ to align them parallelly. If the horizontal distance $d\leq 1.2$ cm and the vertical overlap length $x$ is greater than 1/7 of the particle length, the two clusters are allowed to adhere with a probability of $p=0.02$. The final vertical overlap length $y$ is determined by the black and blue lines in (a).
V.4 Numerical simulations
@@
We employ Monte Carlo simulations to investigate the self-assembly of two-dimensional PLA particles as the overall area shrinks. These simulations adhere to the standard hard particle Monte Carlo approach, with the inclusion of an adhesion mechanism to replicate attractive interactions, commonly known as the Cheerios effect Vella and Mahadevan (2005); Protière (2023). The particles are modeled as two-dimensional rectangles, and each Monte Carlo step involves $N$ trial moves, encompassing translations and rotations of the particles Haji-Akbari et al. (2009); Wan and Glotzer (2021). We set the length of the long side of the particle as the unit length in simulations. Translation steps are fixed at 0.063, approximately 2.2 mm, while rotation steps are 0.033 radians, roughly 1.9 degrees. Overlap checks between particles are conducted using the hard particle Monte Carlo module in HOOMD-blue Anderson et al. (2016), and moves resulting in overlap or boundary intersection are rejected.
@@
To account for the attraction between particles, we incorporate an adhesion mechanism. For simplicity, we consider adhesion only between the long sides of particles and attraction between two particles. Specifically, when the outermost particles of two adjacent clusters (or individual particles) form an angle of $\alpha\leq 5^{\circ}$, we rotate each cluster by $\alpha/2$ to align them in parallel [see an illustration in Fig. 6(b)]. If the horizontal distance $d\leq 1.2$ cm and the vertical overlap length $x$ exceeds 5 mm (1/7 of the particle length), we allow the two clusters to adhere to each other with a probability of $p=0.02$. The final vertical overlap length $y$ is determined by the black line fitting in Fig. 6(a) if $y$ is less than the particle length; otherwise, $y$ is set to the particle length, as indicated by the blue line in the figure.
@@
V.5 Identification of clusters
@@
Acknowledgements.
This work was supported by the National Natural Science Foundation of China (Grant No. 12274330) and the Knowledge Innovation Program of Wuhan-Shuguang (Grant No. 2022010801020125). D.W. also acknowledges the “Xiaomi Young Scholar Program” at Wuhan University. This research was supported in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP).

### v3

@@ IV Conclusion
Inspired by the observation of fallen pine needles self-assembling into clusters in small puddles after rainfall, we conducted an experimental investigation of the dynamic evolution of synthetic particles as the liquid surface itself undergoes shrinkage. We observe that particles tend to adhere to one another, forming distinct clusters as the surface area reduces. We identify the emergence of clusters and systematically explore how cluster sizes and the number of clusters evolve as a function of packing fraction. We also analyze particle rotation during the process and discuss the formation of the final configuration comprising clusters of various orientations. Complementary numerical simulations that capture the two main factors, the adhesion between particles and the reduction of surface area, demonstrate qualitative agreement with our experimental findings. Our study contributes to a better understanding of the self-assembly phenomena occurring when macroscopic particles interact on a liquid surface with a dynamically evolving area.
Acknowledgements.
We thank Prof. Cristina Marchetti for a thorough review of the manuscript and valuable suggestions. This work was supported by the National Natural Science Foundation of China (Grant No. 12274330) and the Knowledge Innovation Program of Wuhan-Shuguang (Grant No. 2022010801020125). D.W. also acknowledges the “Xiaomi Young Scholar Program” at Wuhan University. This research was supported in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP).
DATA AVAILABILITY
The data that support the findings of this article are openly available lx2024.
Appendix A Structure and floating behavior of a particle
Figure 5 illustrates the particle structure. Each particle is 35 mm × 7 mm × 7 mm, with a centrally located groove (20 mm × 4 mm × 1 mm) on each of its four major faces. The particles have a 15% filling ratio, enhancing their buoyancy. Note that the colors in Fig. 5 are for illustration purposes; the actual particles are white, with grooves marked in red for image analysis.
To visualize the meniscus deformation around a particle, we prepared a transparent bowl filled with water and topped with a layer of oil. We placed a particle near the wall of the bowl, with its smaller side facing the wall. We then captured an image of the smaller side from outside the bowl wall. Figure  6(a) shows the image of the floating particle. In Fig. 6(b), we show the estimated contact angle $\theta$, which was measured to be approximately 72°. The deformation of the surface was fitted using the following equation Mahadevan2005:
${y=}{\lambda}{\cot(\theta)}\cdot{\exp\left(\frac{-(x-x_{0})}{\lambda}\right)+}{h_{0}}.$ (1)
Substituting $x_{0}=13.7$ $m$, $h_{0}=11.4$ mm, which were measured from the photograph, and $\theta=72^{\circ}$, the fitting yields a capillary length of $\lambda=1.82$ mm. The fitted curve is shown as the red line in Fig. 6(b).
@@ Appendix A Structure and floating behavior of a particle
Figure 6: Meniscus deformation around a single particle floating on the oil surface. (a) Original photograph of the floating particle. (b) Photograph showing the coordinate system, the estimated contact angle $\theta$, and the fitted curve (red line) representing the surface deformation.
Appendix B Two-particle attraction
@@
To investigate the attraction between two particles, we align them parallel to each other with their large faces facing each other. Maintaining a horizontal separation of $d=1.2$ cm, we systematically vary the initial vertical overlap length $x$ from 1/7 to 6/7 of the particle length, in increments of 1/7 as equal divisions. Upon releasing the particles, they adhere to each other, and we measure the final overlap length $y$ for each initial overlap length $x$, as depicted in Fig. 7(a). Note that the particles do not necessarily remain parallel during the adhesion process, particularly when they are closer. Often, one end of a particle adheres to the other particle first, followed by the rest of the particle adhering subsequently. The black line represents the linear fit of the first five experimental data points, where the final overlap $y$ values are less than the particle length. The blue line denotes the particle length.
@@
Figure 7: Adhesion of two particles. (a) Final overlap length $y$, as a function of initial overlap length $x$, when two particles are released at a fixed horizontal distance $d=1.2$ cm. The inset illustrates the initial and final states. The black line represents the linear fit of the experimental data based on the first five data points, while the blue line denotes the particle length. (b) Determination of the adhesion of two clusters in simulations. When the outermost particles (marked in blue) of two adjacent clusters form an angle $\alpha\leq 5^{\circ}$, we rotate each cluster by $\alpha/2$ to align them in parallel. If the horizontal distance $d\leq 1.2$ cm and the vertical overlap length $x$ is greater than 1/7 of the particle length, the two clusters are allowed to adhere with a probability of $p=0.02$. The final vertical overlap length $y$ is determined by the black and blue lines in (a).
Appendix C Simulation details
@@
We employ Monte Carlo simulations to investigate the self-assembly of two-dimensional PLA particles as the overall area shrinks. These simulations adhere to the standard hard particle Monte Carlo approach, with the inclusion of an adhesion mechanism to replicate attractive interactions, commonly known as the Cheerios effect Mahadevan2005; Protiere2023. The particles are modeled as two-dimensional rectangles, and each Monte Carlo step involves $N$ trial moves, encompassing translations and rotations of the particles Haji-Akbari2009; Wan2021. We set the length of the long side of the particle as the unit length in simulations. Translation steps are fixed at 0.063, approximately 2.2 mm, while rotation steps are 0.033 radians, roughly 1.9 degrees. Overlap checks between particles are conducted using the hard particle Monte Carlo module in HOOMD-blue Anderson2016, and moves resulting in overlap or boundary intersection are rejected.
@@
To account for the attraction between particles, we incorporate an adhesion mechanism. For simplicity, we consider adhesion only between the long sides of particles and attraction between two particles. Specifically, when the outermost particles of two adjacent clusters (or individual particles) form an angle of $\alpha\leq 5^{\circ}$, we rotate each cluster by $\alpha/2$ to align them in parallel [see an illustration in Fig. 7(b)]. If the horizontal distance $d\leq 1.2$ cm and the vertical overlap length $x$ exceeds 5 mm (1/7 of the particle length), we allow the two clusters to adhere to each other with a probability of $p=0.02$. The value of $p$ is chosen to match experimental data. As shown in Fig. 8, varying $p$ influences the average cluster size. This figure presents simulations starting from the same initial configuration but with different $p$ values, while all other parameters remain unchanged. The final vertical overlap length $y$ is determined by the black line fitting in Fig. 7(a) if $y$ is less than the particle length; otherwise, $y$ is set to the particle length, as indicated by the blue line in the figure.
@@
Figure 8: The average cluster size $\langle n_{c}\rangle$ as a function of packaging density $\phi$ for various $p$ values.
Appendix D Identification of clusters
@@
Appendix E Experiments with particles of various densities
We placed two particles initially parallel on the corn oil surface in a head-to-head, end-to-end configuration, meaning the initial overlap length $x$ equals the particle length and the horizontal distance $d$ represents the center-to-center separation. We investigated the attraction time—the time required for two particles to fully adhere—using particles with three different densities: $\rho_{1}=0.65~\text{g}/\text{cm}^{3}$, $\rho_{2}=0.75~\text{g}/\text{cm}^{3}$, and $\rho_{3}=0.86~\text{g}/\text{cm}^{3}$. The particles with $\rho_{2}=0.75~\text{g}/\text{cm}^{3}$ correspond to those described in the main text. Table 1 summarizes the attraction time for various initial separations $d$. For the particles with $\rho_{2}=0.75\,\text{g}/\text{cm}^{3}$, when the initial separation is 2.7 cm, their behavior differs from that observed at smaller initial separations. In this case, their trajectories are easily perturbed by noise after prolonged waiting times, and no reliable trajectory is observed over repeated attempts. Therefore, particles with initial separations of 2.7 cm or greater are not expected to come into contact within the experimental timescale unless additional factors, such as surface area reduction, are considered. When comparing the attraction times of particles with different densities, we observe that lighter particles consistently require less time to come into contact.
Table 1: Table 1. Attraction time (in minutes) for two particles initially separated by a distance $d$.
Initial distance $d$ (cm) Particle density $(\text{g}/\text{cm}^{3})$ $\rho_{1}=0.65$ $\rho_{2}=0.75$ (main text) $\rho_{3}=0.86$
1.2 1 2 5
1.7 2 5 10
2.2 6 24 30
2.7 21 No traj. No traj.
Figure 9: The experimental average cluster size $n_{c}>$ as a function of packing density for three particle densities ($\rho_{1}=0.65\text{g}/\text{cm}^{3}$, $\rho_{2}=0.75\text{g}/\text{cm}^{3}$, $\rho_{3}=0.86\text{g}/\text{cm}^{3}$). The red circles indicate data referenced in the main text.
Figure  9 shows clustering experiments for particles of three different densities. The results indicate a similar trend, where the average cluster size increases with packing density. However, due to differences in initial configurations and the limited number of experimental attempts, we cannot draw definitive conclusions about the influence of particle density on clustering behavior.